# On Selection of Cross-Section Averages in Non-stationary Environments

### v2

@@ 6.1 Auxiliary Results
$=\epsilon^{-2}\frac{1}{T^{1+\tau}}\left(\mathrm{tr}\left[\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\left(\mathbf{f}_{t}^{\prime}\mathbf{f}_{t}\right)\mathbb{E}\left(\overline{\mathbf{e}}_{t}^{\prime}\overline{\mathbf{e}}_{s}\right)\right]\right)$
@@
$\left\|\frac{1}{T^{1+\tau}}\mathbf{F}^{\prime}\mathbf{U}_{i}\right\|=O_{p}(T^{-(1+\tau)/2})$
@@
$=\epsilon^{-2}\frac{1}{T^{1+\tau}}\left(\mathrm{tr}\left[\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\left(\mathbf{f}_{t}^{\prime}\mathbf{f}_{t}\right)\mathbb{E}\left(\mathbf{e}_{i,t}^{\prime}\mathbf{e}_{i,s}\right)\right]\right)$
@@ 6.2 Proof of Proposition 1
$=\underbrace{\ln\mathrm{det}\left(\mathbf{I}_{k}+\left[\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}\right]\overline{\mathbf{Q}}_{M_{0}}^{-1}\right)}_{\mathcal{V}_{DVS}-\mathcal{V}_{DVS}^{0}}+k(g-m)p_{N,T},$ (A.8)
where we used $\ln(a)-\ln(b)=\ln\left(\frac{a}{b}\right)$ and $\mathrm{det}(\mathbf{A}\mathbf{B})=\mathrm{det}(\mathbf{A})\mathrm{det}(\mathbf{B})$. This implies that we will examine the asymptotic behavior of $\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}$. Its behavior is determined by whether $M_{0}\subset M$ (over-specification) or $M\subset M_{0}$ (under-specification). In the latter case, under-specification happens when $M\subset M_{0}$, $M_{0}\cap M\neq\emptyset$ but neither is a weak subset of each other, and when $M\cap M_{0}=\emptyset$. We only analyze the case of $M\subset M_{0}$ as in Margaritella and Westerlund, (2023) or De Vos and Stauskas, (2024), because other cases will lead to the same asymptotic conclusions.
6.3 Part $M\subset M_{0}$
@@
$\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}$ $=\left[\overline{\mathbf{Q}}_{M}-\widehat{\mathbf{Q}}_{M,\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}}\right]-\left[\overline{\mathbf{Q}}_{M_{0}}-\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}\right]+\left[\widehat{\mathbf{Q}}_{M,\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}}-\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}\right]$
@@
$\mathbf{I}$ $=\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{X}_{i}$
$=\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{V}_{i}+\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}$
$+\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}+\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{V}_{i}$
@@
without further restrictions (but negligible under $TN^{-1}=O(1)$. Also,
@@
$=O_{p}(N^{-1})+O_{p}(N^{-1/2}T^{(\tau-1)/2})=o_{p}(1)$ (A.14)
without further restrictions as $\tau\in(0,1)$. Both of these results will be important when determining the orders of many terms.
@@
$\left(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}\right)^{+}=\left(\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\right)^{+}+O_{p}(\xi_{N,T,\tau}^{-1}).$ (A.15)
@@
$\left\|\mathbf{I}_{a}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{V}_{i}\right\|$
$\leq\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{V}_{i}\right\|$
@@
$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|$
@@
$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{V}_{i}\right\|$
@@
$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\left[(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right]T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|$
@@
$\leq\underbrace{T^{-\tau}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-1}\overline{\mathbf{V}}^{\prime}\mathbf{V}_{i}\right\|^{2}\left\|\mathbf{q}_{M}\right\|^{2}}_{O_{p}(N^{-1}T^{-\tau})+O_{p}(N^{-1/2}T^{-(2\tau+1)/2})}$
@@
$+2\underbrace{\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|\mathbf{q}_{M}\right\|^{2}\left\|T^{-1}\overline{\mathbf{V}}^{\prime}\mathbf{V}_{i}\right\|\left\|\overline{\boldsymbol{\Gamma}}\right\|\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|}_{O_{p}(N^{-1}T^{-(1+\tau)/2})+O_{p}(N^{-1/2}T^{-(2+\tau)/2})}$
@@
$+\underbrace{T^{-1}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|\mathbf{q}_{M}\right\|^{2}\left\|\overline{\boldsymbol{\Gamma}}\right\|^{2}\left\|T^{-(1+\tau)/2}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|^{2}}_{O_{p}(T^{-1}\xi_{N,T,\tau}^{-1})}$
$=O_{p}(N^{-1}T^{-\tau})+O_{p}(N^{-1/2}T^{-(2\tau+1)/2}),$ (A.17)
which is the dominating order. Next, we have
@@
$\left\|\mathbf{I}_{b}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
@@
$\leq\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
@@
$+\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$

### v3

@@ 5.2 Auxiliary Results
$=\epsilon^{-2}\frac{1}{T^{1+\tau}}\left(\mathrm{tr}\left[\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\left(\mathbf{f}_{s}^{\prime}\mathbf{f}_{t}\right)\mathbb{E}\left(\overline{\mathbf{e}}_{t}^{\prime}\overline{\mathbf{e}}_{s}\right)\right]\right)$
$=\epsilon^{-2}\frac{1}{T^{1+\tau}}\left(\mathrm{tr}\left[\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\left(\mathbf{f}_{t}^{\prime}\mathbf{f}_{t}\right)\mathbb{E}\left(\mathbb{E}\left(\overline{\mathbf{e}}_{t}^{\prime}\overline{\mathbf{e}}_{s}|\mathcal{F}_{(t-1)\lor(s-1)}\right)\right)\right]\right)$
@@
$\left\|\frac{1}{T^{1+\tau}}\mathbf{F}^{\prime}\mathbf{E}_{i}\right\|=O_{p}(T^{-(1+\tau)/2})$
@@
$=\epsilon^{-2}\frac{1}{T^{1+\tau}}\left(\mathrm{tr}\left[\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\left(\mathbf{f}_{s}^{\prime}\mathbf{f}_{t}\right)\mathbb{E}\left(\mathbf{e}_{i,t}^{\prime}\mathbf{e}_{i,s}\right)\right]\right)$
$=\epsilon^{-2}\frac{1}{T^{1+\tau}}\left(\mathrm{tr}\left[\frac{1}{T^{1+\tau}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\left(\mathbf{f}_{t}^{\prime}\mathbf{f}_{t}\right)\mathbb{E}\left(\mathbb{E}\left(\mathbf{e}_{i,t}^{\prime}\mathbf{e}_{i,s}|\mathcal{F}_{(t-1)\lor(s-1)}\right)\right)\right]\right)$
@@ 5.3 Proof of Proposition 1
$=\ln\mathrm{det}\left(\mathbf{I}_{k}+\left[\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}\right]\overline{\mathbf{Q}}_{M_{0}}^{-1}\right)+k(g-m)p_{N,T}$
$=\ln\mathrm{det}\left(\mathbf{I}_{k}+T^{\tau}\left[T^{-\tau}\left(\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}\right)\right]\overline{\mathbf{Q}}_{M_{0}}^{-1}\right)+k(g-m)p_{N,T},$ (A.8)
where we used $\ln(a)-\ln(b)=\ln\left(\frac{a}{b}\right)$ and $\mathrm{det}(\mathbf{A}\mathbf{B})=\mathrm{det}(\mathbf{A})\mathrm{det}(\mathbf{B})$. This implies that we will examine the asymptotic behavior of $T^{-\tau}\left(\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}\right)$. Its behavior is determined by whether $M_{0}\subset M$ (over-specification) or $M\subset M_{0}$ (under-specification). In the latter case, under-specification happens when $M\subset M_{0}$, $M_{0}\cap M\neq\emptyset$ but neither is a weak subset of each other, and when $M\cap M_{0}=\emptyset$. We only analyze the case of $M\subset M_{0}$ as in Margaritella and Westerlund, (2023) or De Vos and Stauskas, (2024), because other cases will lead to the same asymptotic conclusions.
5.3.1 Case of $M\subset M_{0}$
@@
$T^{-\tau}\left(\overline{\mathbf{Q}}_{M}-\overline{\mathbf{Q}}_{M_{0}}\right)$ $=T^{-\tau}\left[\overline{\mathbf{Q}}_{M}-\widehat{\mathbf{Q}}_{M,\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}}\right]-T^{-\tau}\left[\overline{\mathbf{Q}}_{M_{0}}-\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}\right]$
$+T^{-\tau}\left[\widehat{\mathbf{Q}}_{M,\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}}-\widehat{\mathbf{Q}}_{M_{0},\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}}}\right]$
@@
$\mathbf{I}$ $=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{X}_{i}$
$=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{V}_{i}+\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}$
$+\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}+\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{V}_{i}$
@@
without further restrictions (but negligible under $TN^{-1}=O(1)$. Also, let
@@
$=O_{p}(N^{-1})+O_{p}(N^{-1/2}T^{(\tau-1)/2})$
$=o_{p}(1))=R_{N,T}$ (A.14)
without further restrictions as $\tau\in(0,1)$. Note that the latter will be the dominant component in the further analysis. Both of these results will be important when determining the orders of many terms.
@@
$\left(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}\right)^{+}=\left(\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\right)^{+}+O_{p}(\xi_{N,T,\tau}^{-1}),$ (A.15)
because
$\left\|\left(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}\right)^{+}-\left(\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\right)^{+}\right\|$
$=\left\|\left(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}\right)^{+}\left(\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}-T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}\right)\left(\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\right)^{+}\right\|$
$\leq\left\|\left(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}\right)^{+}\right\|\left\|T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}-\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\right\|\left\|\left(\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\right)^{+}\right\|$
$=O_{p}(\xi_{N,T,\tau}^{-1}).$ (A.16)
@@
$\left\|\mathbf{I}_{a}\right\|$ $=\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{V}_{i}\right\|$
$\leq\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{V}_{i}\right\|$
@@
$+\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|$
@@
$+\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{V}_{i}\right\|$
@@
$+\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\left[(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right]T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|$
@@
$\leq\underbrace{T^{-2\tau}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-1}\overline{\mathbf{V}}^{\prime}\mathbf{V}_{i}\right\|^{2}\left\|\mathbf{q}_{M}\right\|^{2}}_{O_{p}(N^{-1}T^{-2\tau})+O_{p}(N^{-1/2}T^{-(4\tau+1)/2})}$
@@
$+2\underbrace{T^{-\tau}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|\mathbf{q}_{M}\right\|^{2}\left\|T^{-1}\overline{\mathbf{V}}^{\prime}\mathbf{V}_{i}\right\|\left\|\overline{\boldsymbol{\Gamma}}\right\|\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|}_{O_{p}(N^{-1}T^{-(1+2\tau)/2})+O_{p}(N^{-1/2}T^{-(2+3\tau)/2})}$
@@
$+\underbrace{T^{-1-\tau}\left\|(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}-(T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M})^{+}\right\|\frac{1}{N}\sum_{i=1}^{N}\left\|\mathbf{q}_{M}\right\|^{2}\left\|\overline{\boldsymbol{\Gamma}}\right\|^{2}\left\|T^{-(1+\tau)/2}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{V}_{i}\right\|^{2}}_{O_{p}(T^{-(1+\tau)}\xi_{N,T,\tau}^{-1})}$
$=o_{p}(\xi_{N,T,\tau}^{-1})=o_{p}(T^{-\tau}),$ (A.18)
where each of the term is of the lower order than $O_{p}(\xi_{N,T,\tau}^{-1})$. Next, we have
@@
$\left\|\mathbf{I}_{b}\right\|$ $=\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M_{\mathbf{x}}}}-\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}})\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
@@
$\leq\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\mathbf{V}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
@@
$+\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\overline{\mathbf{V}}\mathbf{q}_{M}(T^{-(1+\tau)}\widehat{\mathbf{F}}_{M_{\mathbf{x}}}^{\prime}\widehat{\mathbf{F}}_{M_{\mathbf{x}}})^{+}T^{-(1+\tau)}\mathbf{q}_{M}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$